# Enabling Efficient Optimal Control of Reconfigurable Battery Systems by Unifying System Modeling and Narrowing Search Space

### v1

@@ Abstract
Reconfigurable battery systems (RBSs) are emerging as a promising solution to improving fault tolerance, charge and thermal balance, energy delivery, etc. To optimize these performance metrics, high-dimensional nonlinear integer programming problems need to be formulated and solved. During this process, it is necessary to address multiple challenges stemming from nonlinear battery characteristics, discrete switch states, dynamic system configurations, as well as the curse of dimensionality inherent in large-scale systems. Thus, we propose a unified modeling framework to accommodate various potential configurations of an RBS and even to cover different RBS designs, significantly facilitating the topology design and optimization problem formulation for RBSs. Moreover, to solve the formulated RBS problems, the search space is tailored to encompass only feasible SSVs, thereby ensuring safe system operation while substantially curtailing search efforts. These proposed methods, focusing on unifying the system modeling and narrowing the search space, lay a solid foundation for effectively formulating and efficiently solving RBS optimal control problems. The accuracy and effectiveness of the proposed methods are demonstrated by simulation and experimental tests.
@@
Reconfigurable battery systems, optimal control, battery system modeling, feasible solution space construction.
@@ I Introduction
In response to the urgent imperative of mitigating greenhouse gas emissions, there has been a notable surge in the adoption of renewable energy sources by, e.g., solar panels and wind turbines [1]. Batteries, serving as a primary energy storage medium, play a pivotal role in both integrating renewable energy into power grids and facilitating its application to transportation electrification [2]. Given the intrinsic electrochemical nature of batteries, multiple batteries need to be connected either in series or in parallel to meet diverse voltage, current, or power demands across various applications. For multi-battery systems, inconsistent battery parameters and unbalanced operating states, such as state of charge (SoC), temperature, and state of health (SoH), result in a range of issues spanning from diminished energy delivery [3], accelerated degradation [4] to potential safety hazards [5]. Addressing these issues arising from unbalanced battery operation involves manipulating the operating currents of relevant batteries. However, in many practical battery systems, implementing individual current control becomes infeasible due to the commonly fixed battery connection configuration.
Aimed at mitigating the limitations imposed by fixed battery connections on system performance control, reconfigurable battery systems (RBSs), employing switches to dynamically reconstruct battery connections, are emerging as a promising solution [6]. Various RBS designs, featuring diverse switch arrays and demonstrating different levels of reconfigurability, have been proposed [7, 8, 9, 10, 11, 12]. For instance, Fig. 1 illustrates some RBS structure designs with two to five switches associated with each battery cell. Through switch operations, the battery connection can be reconfigured, enabling the recombination of battery terminal voltages and redistribution of currents among batteries. Then, by designing and implementing appropriate control algorithms, we can potentially improve the RBS’s operation performance in terms of enhanced fault tolerance [13], extended energy delivery [14], balanced battery states [15] and thermal distribution [16], and lower power efficiency [17]. Moreover, RBSs can accommodate batteries of different age, type and even chemistry [18], and facilitate the diagnosis, repair, and replacement of local batteries as well.
@@
In order to fully leverage these potential benefits of RBSs, we need to formulate one specific optimal control problem, typically consisting of an objective function, a system plant, and some necessary constraints, as will be detailed in Section II. The system plant describing the system state evolution is formulated based on the electrical interconnection of all batteries, i.e., the system configuration. Thus, the system configuration, which can be specified by a switch state vector (SSV) describing the on-off states of all switches involved in the RBS, directly influences the evolution of system states and the value of objective function. Motivated by this, the system configuration is selected as one control variable. Through appropriate system reconfiguration, various system performance metrics defined in the objective function can get optimized.
@@
When formulating and solving an RBS optimal control problem, a critical challenge is how to construct a system plant to model an RBS’s dynamically varying system configuration, which has been rarely investigated in the literature. As the control variable, the RBS’s configuration could vary within a wide range. The optimal system configuration is typically approached through an iterative evaluation-and-update process. During each iteration, a batch of configurations (or configuration series if multiple time steps are involved) are first evaluated and then updated according to certain principles. To evaluate the performance of a specific configuration, we need to construct its corresponding system plant, also called the system model, based on all battery cells’ mathematical models as well as equations characterizing their electric connection according to Kirchhoff’s current law (KCL) or voltage law (KVL) [19]. When evaluating another system configuration of the RBS, in which some battery units quit or join the operation and certain local series or parallel connections might change as well, it becomes difficult to directly revise the original system plant since some entry values and even the dimensions of system matrices might change. In this case, it is commonly necessary to construct a new system plant for the different system configuration. However, it is difficult to generate new plants automatically by programming codes due to involving re-indexing the batteries put into operation, rewriting their KCL and KVL equations, and reformulating the system matrices accordingly. Consequently, evaluating various system configurations by manually constructing different plants becomes very inefficient.
@@
To address the issues caused by dynamic reconfiguration, a unified RBS modeling framework is proposed in this work. The formulated system plant, explicitly expressed in terms of the system configuration or SSVs, is capable of covering all possible configurations of an RBS without reconstruction, significantly reducing the computational efforts of evaluating various system configurations. In addition, thanks to the unified explicit expressions of the system plant, constraints for the system state transition can be formulated, allowing various commercial programming solvers to be applied directly.
@@
When solving an RBS optimal control problem, another critical challenge is how to efficiently search the optimal solution. During the process of iterative evaluation and update, the number of possible system configurations experiences exponential growth with the number of switches involved in the RBS as well as the number of time steps within the control horizon, as will be demonstrated in Section IV. Thus, for large-scale RBSs, it becomes very time-consuming to search the optimal solution among a huge number of possible system configurations.
@@
To alleviate the influence of the “curse of dimensionality”, we propose to actively construct the feasible solution space. By filtering all SSVs incurring open circuits, short circuits, undesired system terminal voltage, or duplicate configurations, the size of the constructed feasible solution space is much smaller than that of the complete solution space exponentially expanding with the number of switches. In addition to ensuring safe operation, the proposed method for narrowing the search space can dramatically enhance the search efficiency when optimizing certain RBS operating performance.
@@
The reminder of this work are organized as follows. First, the general optimal control problem for RBSs will be formulated in Section II. In Section III, the unified modeling framework for RBSs will be introduced along with simulation and experimental tests to demonstrate the modeling accuracy. Then, the methodology to narrow the search space will be investigated in Section IV, and its performance will be compared with conventional methods. Finally, conclusions will be presented in Section V.
II Optimal control problem formulation for RBSs
@@
$min~{}J(X_{0},U)$
$s.t.$
$X_{k}=\Phi(u_{k},X_{k-1}),k=1,2,\ldots,K,$
$g_{l}(X_{0},U)\geq 0,l=1,2,\ldots,L,$
$h_{r}(X_{0},U)=0,r=1,2,\ldots,R,$
@@
In the objective function $J$, $X_{k}$ denotes the system state at $t=k\Delta t$, and $U=[u_{1},u_{2},\ldots,u_{K}]$ the input trajectory, in which $\Delta t$ is the duration of each time step, and $K$ is the number of time steps considered during the optimization process. Among the constraints, the first one describes the system plant, followed by some inequality and equality constraints, and $\mathcal{U}^{K}$ denotes the constraint set for the input trajectory.
Different from traditional battery systems, RBS enables the possibility to select the system configuration described by switch state vectors (SSVs) as one control variable or system input for the optimal control. Specifically, for the $k$-th time step $((k-1)\Delta t,k\Delta t]$, the corresponding input $u_{k}$ is expressed as follows based on the on-off state of each switch, denoted by $s_{n}\in\mathcal{B}=\{0,1\}$, $n=1,2,\ldots,N_{S}$, where $N_{S}$ is the total number of switches in the RBS.
$u_{k}=[s_{1},s_{2},\ldots,s_{N_{S}}]\in\mathcal{B}^{N_{S}}.$ (1)
The general process of iterative evaluation and update of SSVs for searching the optimal solution is illustrated in Fig. 2, in which the right part describes the performance evaluation of SSVs, the left part presents the search space of SSVs, and the optimization algorithms at the bottom generates the updated SSVs for each iteration. Among these three parts, the first two parts lay the foundation for adopting various optimization algorithms. Thus, this work will be mainly focused on boost the efficiency of evaluating and searching SSVs.
Figure 2: Structural diagram of proposed framework.
For efficiently evaluating the RBS performance metric described by the objective function $J_{k}$, in Section III we will develop a unified system plant with the SSV as input to accommodate diverse system configurations, corresponding to the unified state transition function $\Phi(SSV,X)$ shown in Fig. 2. To further improve the searching efficiency, as shown in Fig. 2, instead of directly searching within the complete search space $\mathcal{B}^{N_{S}}$, we will identify only those feasible SSVs to form a dramatically narrowed search space in Section IV.
@@ III Unified modeling framework for RBSs
Different structures can be designed for RBSs by changing the number of switches around each battery cell and the cell-switch connection pattern. For instance, as shown in Fig. 1, the RBS design (a) can encompass other designs (b)–(f) by selectively removing specific switch or wire branches (i.e., disconnecting the corresponding switches). Thus, we will take RBS design (a) as an example to introduce the unified modeling framework for RBSs.
III-A Describing battery reconfiguration by switch resistance variables
To model an RBS, not only all components, including batteries, switches, wires, contactors, etc., but also the system configuration, i.e., how all components are interconnected with each other, needs to be described. The diagram for design (a) is presented in Fig. 3, in which the $n$-th battery cell, $n=1,2,\ldots,N_{B}$, is represented by a second-order equivalent circuit model (ECM) [20], and $N_{B}$ is the total number of battery cells. In this battery cell model, $v_{n}$ denotes the open circuit voltage (OCV), $R_{0,n}$ denotes the internal resistance, and two parallel resistor-capacitor (RC) pairs are used for describing the influence of the lithium diffusion process on the cell terminal voltage. Other types of battery models can be deployed as well if necessary. Around the $n$-th battery cell, the five switches are labeled as $S^{m}_{n}$, $m=1,2,\ldots,5$, and each switch’s state is denoted by $s^{m}_{n}\in\{0,1\}$. If a switch’s state is on, i.e., conductive, $s^{m}_{n}=1$ and its resistance is represented by its static drain to source on resistance $R_{\text{ds(on)},n}^{m}$. Otherwise, $s^{m}_{n}=0$ and the branch with an off-state switch can be removed.
@@ III-A Describing battery reconfiguration by switch resistance variables
Thus, the system configuration is not fixed but depends on the on or off switch states. When the RBS is dynamically reconfigured by switch operation, i.e., adjusting the states of certain switches, some local meshes in the diagram change accordingly, as shown in Fig. 3. Consequently, we have to reconstruct the system model since some battery cells need to be re-indexed, and their corresponding KCL or KVL equations need to be rewritten as well.
To boost the modeling efficiency of RBSs by avoiding frequent system model reconstruction, it is desired to keep all meshes unchanged despite the dynamic battery reconfiguration. Thus, instead of directly removing those off-state switch branches, all branches are reserved to form the same meshes during reconfiguration. Further, to reflect the changes in switch states, different resistances are assigned to switches with on and off states, and they are combined into a switch resistance variable $R^{S_{m}}_{n}$ through the binary switch state $s^{m}_{n}$, as proposed in (2). In this expression, $R_{\text{ds(off)},n}$ denotes the switch’s drain to source off resistance, which is typically at $M\Omega$ level for practical MOSFETs [21]. Moreover, to simplify the denotation, the resistance of the connection wire along the switch branch, denoted by $R_{\text{wire},n}^{m}$, is incorporated into $R^{S_{m}}_{n}$.
@@
According to (2), the switch branch’s resistance $R^{S_{m}}_{n}=R_{\text{\text{ds(on)}},n}^{m}+R_{\text{wire},n}^{m}$ if the switch is on, or $R^{S_{m}}_{n}=R_{\text{\text{ds(off)}},n}^{m}+R_{\text{wire},n}^{m}$ if the switch is off. Then, any switch state change corresponding to the dynamic reconfiguration can be reflected by the finite value of the switch resistance variable. Through this way, those off-state switch branches are reserved, and we can always maintain a full system configuration (FSC) involving all switch, battery, and wire branches. Once an SSV is specified by all switch states, all switch resistances will be determined based on (2), and the corresponding system configuration FSC(SSV) will be obtained. Most importantly, the FSC features stable meshes regardless of the reconfiguration, significantly facilitating the formulation of KCL and KVL equations. Thanks to this property of the FSC, it becomes possible to develop a unified system plant to cover all system configurations of an RBS.
III-B Formulation of the unified RBS model
To formulate the state-space representation for the RBS plant or model based on the system’s FSC, the states of battery cells shown in Fig. 3, including each cell’s the resistor currents $i_{n}^{R_{1}}$ and $i_{n}^{R_{2}}$ in the RC pairs and the OCV $v_{n}$, are chosen to compose the RBS’s state vector as follows.

### v2

# Fundamental Techniques for Optimal Control of Reconfigurable Battery Systems: System Modeling and Feasible Search Space Construction

@@ Abstract
Reconfigurable battery systems (RBSs) are emerging as a promising solution to improving fault tolerance, charge and thermal balance, energy delivery, etc. To optimize these performance metrics of RBSs, high-dimensional nonlinear integer programming problems need to be formulated and solved. To accomplish this, it is necessary to address several critical challenges stemming from nonlinear battery characteristics, discrete switch states, dynamic system configurations, as well as the curse of dimensionality inherent in large-scale RBSs. Thus, we propose a unified modeling framework to accommodate various possible configurations of an RBS and even to cover different RBS designs and their hybrid combinations, enabling the problem formulation for the RBS optimal control and facilitating the RBS topology design. Further, to solve the formulated RBS optimal control problems, the search space is narrowed to encompass only the feasible solutions, thereby ensuring safe battery connections while substantially curtailing search efforts. These proposed techniques, focusing on unifying the system modeling and narrowing the search space, lay a solid foundation for effectively formulating and efficiently solving RBS optimal control problems. The accuracy and effectiveness of the proposed techniques are demonstrated by both simulation and experimental tests.
@@
Reconfigurable battery systems, optimal control, battery system modeling, feasible search space.
@@ I Introduction
In response to the urgent imperative of mitigating greenhouse gas emissions, there has been a notable surge in the adoption of renewable energy sources, such as solar panels and wind turbines [1]. Batteries, serving as a primary energy storage medium, play a pivotal role in both integrating renewable energy into power grids and facilitating its application to transportation electrification [2]. Given the intrinsic electrochemical nature of batteries, multiple batteries need to be connected either in series or in parallel to meet diverse voltage, current, or power demands across various applications. For multi-battery systems, inconsistent battery parameters and unbalanced operating states, such as state of charge (SoC), temperature, and state of health (SoH), result in a range of issues spanning from diminished energy delivery [3], accelerated degradation [4] to potential safety hazards [5]. Addressing these issues arising from unbalanced battery operation involves manipulating the operating currents of relevant batteries. However, in many practical battery systems, implementing current control of each individual cell is infeasible due to the commonly fixed battery connection configuration.
Aimed at mitigating the limitations imposed by fixed battery connections on system performance control, reconfigurable battery systems (RBSs), employing switches to dynamically reconstruct battery connections, are emerging as a promising solution [6]. Various RBS designs, featuring diverse switch arrays and demonstrating different levels of reconfigurability, have been proposed [7, 8, 9, 10, 11, 12]. For instance, Fig. 1 illustrates some RBS structure designs with two to five switches associated with each battery cell. Through switch operations, the battery connection can be reconfigured, enabling the recombination of battery terminal voltages and redistribution of currents among batteries. Then, by designing and implementing appropriate control algorithms, one can potentially improve the RBS’s operation performance in terms of enhanced fault tolerance [13], extended energy delivery [14], balanced battery states [15] and thermal distribution [16], and greater power efficiency [17]. Moreover, RBSs can accommodate batteries of heterogeneous age, type and even chemistry [18], and facilitate the diagnosis, repair, and replacement of local cells as well.
@@
In order to fully leverage these potential benefits of RBSs, we need to formulate one specific optimal control problem, typically consisting of an objective function, a system model, and some necessary constraints, as will be detailed in Section II. The system model describing the system state evolution is formulated based on the electrical interconnection of all batteries, i.e., the system configuration. The system configuration of an RBS can be specified by a switch state vector (SSV) composed of the on-off states of all switches involved in the system. The evolution of system states and the value of objective function are directly influenced by the specified system configuration. Motivated by this, the RBS’s configuration can be selected as one control variable. Through appropriate reconfiguration, various system performance metrics defined in the objective function are expected to be optimized.
@@
To formulate and solve an optimal control problem for the RBS, a critical challenge is how to construct a system model to simulate the dynamically varying system configuration, which has been rarely investigated in the literature. As the control variable, the RBS’s configuration could vary within a wide range. The optimal system configuration is typically approached through an iterative evaluation-and-update process. During each iteration, a batch of configurations (or configuration series if multiple decision steps are involved) are first evaluated and then updated according to certain principles. To evaluate the performance of a specific configuration, we need to construct its corresponding system model based on all battery cells’ mathematical models as well as equations characterizing their electric connection according to Kirchhoff’s current law (KCL) or voltage law (KVL) [19]. When evaluating another system configuration of the RBS, in which some battery cells quit or join the operation and certain local series or parallel connections might change as well, it becomes difficult to directly revise the original system model since some entry values and even the dimensions of system matrices might change. In this case, it is commonly necessary to construct a new system model. However, it is challenging to generate new models automatically by programming codes due to involving re-indexing the batteries put into operation, rewriting their KCL and KVL equations, and reformulating the system matrices accordingly. Consequently, manually constructing different models for evaluating various system configurations becomes very inefficient.
@@
To address the issues caused by dynamic reconfiguration, we propose the concept of complete system configuration (CSC) to unify the modeling of various RBS configurations in Section III. The formulated system model based on the CSC, explicitly expressed in terms of the system configuration or SSVs, is capable of covering all possible configurations of an RBS without reconstruction, significantly reducing the computational efforts of evaluating various system configurations. Thanks to this modeling approach, a single model constructed for design (a) in Fig. 1 can represent all other designs shown in the figure, as well as hybrid combinations of these designs.
@@
When solving an RBS optimal control problem, another critical challenge is how to efficiently search the optimal solution. During the process of iterative evaluation and update, the number of possible system configurations experiences exponential growth with the number of switches involved in the RBS as well as the number of decision steps within the decision horizon, as will be demonstrated in Section IV. Thus, for large-scale RBSs, it becomes very time-consuming to search for the optimal solution among a huge number of possible system configurations.
@@
To alleviate the influence of the “curse of dimensionality”, we will propose an approach to actively construct the feasible solution space in Section IV. By filtering all SSVs incurring open circuits, short circuits, undesired system voltages, or duplicate configurations, the size of the constructed feasible solution space is much smaller than that of the complete solution space exponentially expanding with the number of switches. In addition to ensuring safe operation, the proposed approach for narrowing the search space can dramatically enhance the search efficiency when solving optimal control problems for RBSs.
@@
The remainder of this work is organized as follows. First, the general optimal control problem for RBSs is formulated in Section II. In Section III, the unified modeling framework for RBSs is introduced along with simulation and experimental tests to demonstrate the modeling accuracy. Then, the approach to narrow the search space is investigated in Section IV, and its performance is compared with conventional methods. Finally, conclusions are presented in Section V.
II Formulation of a general optimal control problem for RBSs
@@
$\min~{}J(X_{0},U,D)$
s.t. $X_{k}=\Phi(X_{k-1},u_{k},D_{k}),\quad k=1,2,\ldots,K,$
$g_{l}(X_{0},U,D)\geq 0,\quad l=1,2,\ldots,L,$
$h_{r}(X_{0},U,D)=0,\quad r=1,2,\ldots,R,$
@@
In the objective function $J$, $X_{k}$ denotes the system state vector at $t=k\Delta t$, $U=[u_{1},u_{2},\ldots,u_{K}]$ the input trajectory, and $D=[D_{1},D_{2},\ldots,D_{K}]$ the disturbances. Notation $\Delta t$ is the duration of each sampling interval and $K$ is the number of sampling intervals considered during the decision horizon. Then, $X_{0}$ is the initial system state vector. Among the constraints, the first one describes the system model, followed by some inequality and equality constraints, and $\mathcal{U}^{K}$ denotes the constraint set for the input trajectory.
Different from a traditional battery system, RBS enables the possibility to select the system configuration, described by switch state vectors (SSVs), as one control variable or system input for optimal control. Specifically, for the $k$-th sampling interval $((k-1)\Delta t,k\Delta t]$, the corresponding input $u_{k}$ is expressed as follows based on the on-off state of each switch, denoted by $s_{n,k}\in\mathcal{B}=\{0,1\}$, $n=1,2,\ldots,N_{S}$, where $N_{S}$ is the total number of switches in the RBS.
$u_{k}=[s_{1,k},s_{2,k},\ldots,s_{N_{S},k}]\in\mathcal{B}^{N_{S}}.$ (1)
Moreover, the system’s load profile, commonly specified in terms of system power or current, is regarded as an external disturbance $D$ in the above problem.
The general process of iterative evaluation and update of SSVs for searching the optimal solution is illustrated in Fig. 2, in which the right part describes the performance evaluation of SSVs, the left part presents the search space of SSVs, and the optimization algorithms at the bottom generates the updated SSVs for each iteration. Among these three parts, the first two parts lay the foundation for adopting various optimization algorithms. Thus, this work is mainly focused on boosting the efficiency of evaluating and searching SSVs.
Figure 2: Structural diagram of the proposed framework.
For efficiently evaluating the RBS performance metric described by the objective function $J$, in Section III, a unified system model with the SSV as input is developed to accommodate diverse system configurations, corresponding to the unified state transition function $\Phi(X,SSV,D)$ shown in Fig. 2. The objective function $J$ depends on the vector of all battery cell currents $I_{B}$, the system’s terminal voltage $v_{t}$, and the system’s state variable vector $X$. To further improve the searching efficiency, instead of directly searching within the complete search space $\mathcal{B}^{N_{S}}$, we will identify only those feasible SSVs to form a dramatically narrowed search space, i.e., the feasible search space, in Section IV.
@@ III Unified modeling framework for RBSs
Different structures can be designed for RBSs by changing the number of switches around each battery cell and the cell-switch connection patterns. For instance, as shown in Fig. 1, RBS design (a) can be used to realize other designs (b)–(f) by selectively removing specific switch or wire branches (i.e., disconnecting the corresponding switches). Thus, we will take RBS design (a) as an example to introduce the unified modeling framework for RBSs.
III-A Complete system configuration for unifying the RBS modeling
To model an RBS, not only all components, such as batteries, switches, wires, and contactors, but also the system configuration, i.e., how all components are interconnected with each other, needs to be described. The diagram for RBS design (a) is presented in Fig. 3, in which the $n$-th battery cell, $n=1,2,\ldots,N_{B}$, is represented by a second-order equivalent circuit model (ECM) [20], and $N_{B}$ is the total number of battery cells. In this battery cell model, $v_{n}$ denotes the open circuit voltage (OCV), $R_{0,n}$ denotes the internal resistance, and two parallel resistor-capacitor (RC) pairs are used for describing the lithium diffusion process. Other types of battery models can be deployed as well if necessary. The five switches associated with the $n$-th battery cell are labeled as $S^{m}_{n}$, $m=1,2,\ldots,5$, and each switch’s state is denoted by $s^{m}_{n}\in\{0,1\}$. If a switch’s state is on, i.e., conducted, $s^{m}_{n}=1$ and its resistance is represented by its static drain-to-source on resistance $R_{\text{ds(on)},n}^{m}$. Otherwise, the off switch state $s^{m}_{n}=0$ indicates that the corresponding branch is disconnected.
@@ III-A Complete system configuration for unifying the RBS modeling
Clearly, an RBS’s configuration depends on the on or off switch states. If those branches including any off-state switches are removed from the diagram in Fig. 3, some local meshes are changed accordingly. Consequently, we need to reconstruct the system model since some battery cells need to be re-indexed, and their corresponding KCL or KVL equations need to be rewritten as well. It is challenging to do so automatically by programming codes. Thus, it becomes very inefficient to model the RBS under dynamic reconfiguration.
To boost the RBS modeling efficiency by avoiding frequent model reconstruction, it is better to keep all meshes unchanged despite the dynamic reconfiguration. Thus, instead of directly removing those off-state switch branches, all branches are reserved so that we can maintain a complete system configuration (CSC) involving all switches, batteries, and wire branches. The CSC features stable meshes regardless of any reconfiguration, significantly facilitating the formulation of KCL and KVL equations. As a result, it becomes possible to develop a unified system model to cover all configurations of an RBS.
When modeling the CSC of an RBS, each switch is characterized by a switch resistance relying on the switch state. Then, the resistance of switch $S^{m}_{n}$ is denoted by $R^{S_{m}}_{n}$ and expressed through the binary switch state $s^{m}_{n}$ as follows.
@@
[…]
III-B Formulation of an RBS model based on its CSC
To formulate the state-space representation for the RBS model based on its CSC, the states of battery cells shown in Fig. 3, including each cell’s resistor currents $i_{n}^{R_{1}}$ and $i_{n}^{R_{2}}$ in the RC pairs and the OCV $v_{n}$, are chosen to compose the RBS state vector as follows.